# Spectroscopic and X-ray Modeling of the Strong Lensing Galaxy Cluster MACS J0138.0-2155

### v1

@@ Abstract
We model the total mass and galactic substructure in the strong lensing galaxy cluster MACS J0138.0-2155 using a combination of Chandra X-ray data, Multi-Unit Spectroscopic Explorer (MUSE) spectroscopy, and Hubble Space Telescope imaging. MACS J0138.0-2155 lenses a source galaxy at $z=1.95$ which hosts two strongly lensed supernovae, Requiem and Encore. We find MACS J0138.0-2155 to have an X-ray temperature of $6.7\pm 0.4$ keV and a velocity dispersion of cluster member galaxies of $718^{+132}_{-182}$ km s-1, which indicate a cluster mass of $\sim 5\times 10^{14}M_{\odot}$. The round morphology of the X-ray emission indicates that this cluster is relaxed with an ellipticity within the lensing region of $e=0.12\pm 0.03$. Using 18 of the brightest, non-blended, quiescent galaxies, we fit the cluster specific Faber-Jackson relation, including a set of 81 variations in the analysis choices to estimate the systematic uncertainties in our results. We find a slope of $\alpha=0.26\pm 0.06(\mathrm{stat.})\pm 0.03(\mathrm{sys.})$ with an intrinsic scatter of $\Delta\sigma=31^{+8}_{-6}(\mathrm{stat.})\pm 4(\mathrm{sys.})$ km s-1 at a reference velocity dispersion of $\sim 220$ km s-1. We also report on significant galaxies along the line-of-sight potentially impacting the lens modeling, including a massive galaxy with stellar velocity dispersion of $291\pm 3$ km s-1 which lies close in projection to the central cluster galaxy. This galaxy is part of a small group at a slightly higher redshift than the cluster.
@@ 2.2 HST Imaging
Galaxy photometry was modeled using the tractor algorithm (Lang et al., 2016). In brief, the “black point” (sky background) and error are measured from several different blank patches across the field, which all give consistent values. The point spread function (PSF) is measured from a nearby star which is modeled in tractor as a sum of three Gaussians. Galaxies in the field were modeled as follows: If they have no other galaxies within 1” of their center, a 1”x1” cutout around the galaxy is taken. If there are multiple galaxies within that distance, a larger cutout including all the nearby galaxies is taken, with at least 0.5” of room on each side. In this case the light profile from each galaxy is fit jointly. Most galaxies are modeled as a sum of an exponential and de Vaucouleur profile (de Vaucouleurs, 1953). For small and faint galaxies, we use an exponential only; of the galaxies in our final sample, this applies to ‘Q’ and ‘R’. Fluxes are converted to magnitudes using the AB mag zero-point computed from the PHOTFLAM and PHOTPLAM keywords in the reduced image and extinction corrected taking the galactic extinction values from the NASA/IPAC Extragalactic Database (NED)  [^1] https://ned.ipac.caltech.edu/Documents/References/
@@ 3.1 Spectroscopic modeling
More detailed spectral modeling was done for the bright, non-blended cluster members using pPXF. This analysis used the available spectra in the rest-frame wavelength range 3850-4500 Å, spanning the Ca H&K and G-band absorption features.  As input stellar templates to pPXF, we use data release 3 of the X-Shooter Library (Verro et al., 2022), which contains spectra of 630 stars. Since the spectral resolution of X-Shooter exceeds that of MUSE, these templates were convolved with a Gaussian line spread function to match the intrinsic resolution of our observed spectra.
@@ 3.2 Faber-Jackson Relation
Firstly, in mapping the Faber-Jackson relation of the cluster, we must make a selection for which member galaxies to include. We select cluster galaxies, with sufficient spectral signal-to-noise to reliably measure a stellar velocity dispersion via absorption features. Ultimately, the 18 galaxies we chose to include in this relation were those which were brighter than $m_{\mathrm{F555W}}=23.5$, within 3,000 km s-1 of the central recessional velocity of the cluster, which did not appear to be spatially blended in the available MUSE data, and had a pPFX fit $\chi^{2}$ less than 2.5 and an error on the stellar velocity dispersion less than 75 km s-1.
@@
where $\sigma$ is a given galaxy’s velocity dispersion, $\sigma_{\mathrm{ref}}$ is a constant for the reference velocity dispersion at a specified magnitude (here taken to be 20), $L$ is the luminosity of the member galaxy, $L_{0}$ is the reference luminosity (corresponding to a magnitude of 20), and $\alpha$ is the power-law slope.
We fit the Faber-Jackson relation following the methodology outlined in Appendix B of Bergamini et al. (2019). Specifically, we employ a Markov Chain Monte Carlo (MCMC) method with emcee’s affine invariant ensemble sampler (Goodman & Weare, 2010; Foreman-Mackey et al., 2013) to find the values of $\sigma_{\mathrm{ref}}$ and $\alpha$ which best fit our population. Additionally, we fit the variable $\Delta\sigma$ representing the intrinsic scatter in the relation. These three variables were fit using a log prior, constrained to the ranges $\alpha\in(0,0.6)$, $\sigma_{ref}\in(50,500)$, and $\Delta\sigma\in(0,200)$. Our log-likelihood function follows Appendix B of Bergamini et al. (2019).
@@
Each of the 81 possible permutations of these variables was run through pPXF with the population of 18 galaxies as described above to derive a new sample of stellar velocity dispersions, and the resulting spread provides us with the systematic error for our measurements. Fits with $\chi^{2}$ greater than 2.5 or an error on the stellar velocity dispersion greater than 75 km s-1 were removed from the sample for that run. We then re-fit the Faber-Jackson relation for each of the 81 realizations giving a range of values for our power-law fit parameters.
@@ 4.1 Faber-Jackson Fit
Galaxy RA DEC F555w MARZ pPXF Velocity
Label Magnitude Redshift Redshift Dispersion (km s-1)
@@ 4.3 Significant Galaxies along the Line-of-Sight
As previously noted, some of the galaxies in our full redshift sample are not cluster members, but still lie along the line of sight to the lensed source and thus affect the lens modeling. Notably, galaxy A is the brightest in the field besides the central galaxy and one of the closest galaxies in projection to the central. Based on the pPXF analysis this galaxy has a velocity offset from the central cluster redshift of 7291 km s-1 and a velocity dispersion of 291 km s-1, corroborating that the galaxy is of high mass. The spectrum and pPFX fit for this galaxy are shown in Figure 7. Along with A, we find four, fainter galaxies with similar redshifts, AF, AK, BC, and BD.
Figure 7: MUSE spectrum of galaxy A (black) overlaid with pPXF’s best fitting model (red) and residuals (green). The spectrum is shown from 3850Å to 4500Å and includes visible Calcium H & K and G-band absorption features. The feature shown in the gray region is an artifact of the sky subtraction.
@@ 4.4 X-ray Properties
From the Chandra analysis, we find that MACS0138 has an $r_{500}$ X-ray temperature of $6.7\pm 0.4$ keV and a soft band (0.5-2 keV) luminosity of $5.5\pm 0.1\times 10^{44}$ ergs s-1. The temperature is consistent for different aperture choices; for example, the $r_{2500}$ temperature is found to be $6.5^{+0.4}_{-0.3}$ keV and the core-cropped $r_{500}$ temperature is found to be $6.6^{+0.9}_{-0.8}$ keV. Based on the $r_{500}$ temperature and using the $M-T$ relation from Mantz et al. (2016), we estimate the mass of the cluster to be $M_{500}=5.2^{+1.5}_{-1.2}\times 10^{14}M_{\odot}$ where the uncertainties include both the X-ray temperature uncertainties and the uncertainties on the fit parameters in Mantz et al. (2016).
@@ 5 Discussion
For the Faber-Jackson relation, Bergamini et al. (2019) present results for three strong lensing clusters at similar redshifts to MACS0138 and find Faber-Jackson slopes of $\alpha=0.28\pm 0.02$, $0.27\pm 0.03$, and $0.27\pm 0.04$. Their results are very consistent with our result of $\alpha=0.26\pm 0.06(\mathrm{stat.})\pm 0.03(\mathrm{sys.})$, though we note that their reference velocity dispersions as slightly higher than ours ($\sim 300$ km s-1). Our measured intrinsic scatter is also consistent with the clusters in Bergamini et al. (2019) with the exception of Abell S1063 for which they measure a higher scatter.
In the final stages of preparation of this paper, the preprint Granata et al. (2024) was posted, which also fits the Faber-Jackson relation in MACS0138. Despite some differences in the two analyses, our results are completely consistent; they find $\alpha=0.25\pm 0.05$ and scatter $\Delta\sigma=25^{+4}_{-6}$ km s-1 at a similar reference velocity dispersion (206 km s-1 compared to our 223 km s-1). Differences in the two analyses include that we use a somewhat larger sample of galaxies (18 vs. 13) reflecting different quality cuts, they use a deeper MUSE cube obtained by their team, and they use F160W magnitudes compared to our use of F [^5] W. While they do investigate the affects of signal-to-noise on their velocity dispersion measurements, they do not do the type of systematics test we employ; our results indicate that their uncertainties may be underestimated. However, the overall consistency of the results of these independent works speaks to their robustness. We cannot compare our velocity dispersion measurements for individual galaxies at this time as their measurements are not included in the preprint version of Granata et al. (2024).

### v2

@@ Abstract
We model the total mass and galactic substructure in the strong lensing galaxy cluster MACS J0138.0-2155 using a combination of Chandra X-ray data, Multi-Unit Spectroscopic Explorer (MUSE) spectroscopy, and Hubble Space Telescope imaging. MACS J0138.0-2155 lenses a source galaxy at $z=1.95$ which hosts two strongly lensed supernovae, Requiem and Encore. We find MACS J0138.0-2155 to have an X-ray temperature of $6.7\pm 0.4$ keV and a velocity dispersion of cluster member galaxies of $718^{+132}_{-182}$ km s-1. These lead to the mass estimates for the cluster of $M_{500}=5.2^{+1.5}_{-1.2}\times 10^{14}M_{\odot}$ from the X-ray results and $M_{200}\approx 3.6^{+2.0}_{-2.7}\times 10^{14}M_{\odot}$ from the velocity dispersion results. The round morphology of the X-ray emission indicates that this cluster is relaxed with an ellipticity within the lensing region of $e=0.12\pm 0.03$. Using 18 of the brightest, non-blended, quiescent galaxies, we fit the cluster specific Faber-Jackson relation, including a set of 81 variations in the analysis choices to estimate the systematic uncertainties in our results. We find a slope of $\alpha=0.26\pm 0.06(\mathrm{stat.})\pm 0.03(\mathrm{sys.})$ with an intrinsic scatter of $\Delta\sigma=31^{+8}_{-6}(\mathrm{stat.})\pm 4(\mathrm{sys.})$ km s-1 at a reference velocity dispersion of $\sim 220$ km s-1. We also report on significant galaxies along the line-of-sight potentially impacting the lens modeling, including a massive galaxy with stellar velocity dispersion of $236\pm 3$ km s-1 which lies close in projection to the central cluster galaxy. This galaxy is part of a small group at a slightly higher redshift than the cluster.
@@ 2.2 HST Imaging
Galaxy photometry was modeled using the tractor algorithm (Lang et al., 2016). In brief, the “black point” (sky background) and error are measured from several different blank patches across the field, which all give consistent values. The point spread function (PSF) is measured from a nearby star which is modeled in tractor as a sum of three Gaussians. Galaxies in the field were modeled as follows: If they have no other galaxies within 1” of their center, a 1”x1” cutout around the galaxy is taken. If there are multiple galaxies within that distance, a larger cutout including all the nearby galaxies is taken, with at least 0.5” of room on each side. In this case the light profile from each galaxy is fit jointly. Most galaxies are modeled as a sum of an exponential and de Vaucouleur profile (de Vaucouleurs, 1953). For small and faint galaxies, we use an exponential only; of the galaxies in our final sample, this applied to galaxies ‘Q’ and ‘R’. Fluxes are converted to magnitudes using the AB mag zero-point computed from the PHOTFLAM and PHOTPLAM keywords in the reduced image and extinction corrected taking the galactic extinction values from the NASA/IPAC Extragalactic Database (NED)  [^1] https://ned.ipac.caltech.edu/Documents/References/
@@ 3.1 Spectroscopic modeling
More detailed spectral modeling was done for the bright, non-blended cluster members using pPXF. This analysis used the available spectra in the rest-frame wavelength range 3850-4500 Å, spanning the Ca H&K and G-band absorption features.  As input stellar templates to pPXF, we use data release 3 of the X-Shooter Library (XSL) (Verro et al., 2022), which contains spectra of 630 stars. In order to prevent over-fitting with low-quality templates, we exclude any XSL spectrum with a mean SNR per pixel lower than 3 in the wavelength range used for spectral fitting. When fitting a spectrum from the MUSE data, pPXF creates a linear combination of template stellar spectra to best match the input galaxy’s spectrum. We first fit a bright member galaxy’s spectrum, and use the template combination provided by pPXF for that galaxy as a fixed “optimal template” for the rest of the cluster members in our sample. Since the spectral resolution of X-Shooter exceeds that of MUSE, these templates were convolved with a Gaussian line spread function to match the intrinsic resolution of our observed spectra.
@@ 3.2 Faber-Jackson Relation
Firstly, in mapping the Faber-Jackson relation of the cluster, we must make a selection for which member galaxies to include. We select cluster galaxies, with sufficient spectral signal-to-noise to reliably measure a stellar velocity dispersion via absorption features. Ultimately, the 18 galaxies we chose to include in this relation were those which were brighter than $m_{\mathrm{F555W}}=23.5$, within 3,000 km s-1 of the central recessional velocity of the cluster, which did not appear to be spatially blended in the available MUSE data, and had a pPFX fit $\chi^{2}$ less than 2.5 and an error on the stellar velocity dispersion less than 75 km s-1. We do not include the central galaxy in the Faber-Jackson relation; cluster central galaxies have a different formation history than satellite galaxies and do not follow the same Faber-Jackson relation (e.g. von der Linden et al., 2007). We separately model the dynamics of the central galaxy and jointly fit these with the lens modeling in a forthcoming paper (O’Donnell et al. in preparation).
@@
where $\sigma$ is a given galaxy’s velocity dispersion, $\sigma_{\mathrm{ref}}$ is a constant for the reference velocity dispersion at a specified magnitude (here taken to be $m_{F555W}=20$), $L$ is the luminosity of the member galaxy, $L_{0}$ is the reference luminosity (corresponding to a magnitude of $m_{F555W}=20$), and $\alpha$ is the power-law slope.
We fit the Faber-Jackson relation following the methodology outlined in Appendix B of Bergamini et al. (2019). Specifically, we employ a Markov Chain Monte Carlo (MCMC) method with emcee’s affine invariant ensemble sampler (Goodman & Weare, 2010; Foreman-Mackey et al., 2013) to find the values of $\sigma_{\mathrm{ref}}$ and $\alpha$ which best fit our population. Additionally, we fit the variable $\Delta\sigma$ representing the intrinsic scatter in the relation. These three variables were fit using a log prior, constrained to the ranges $\alpha\in(0,0.6)$ km s-1, $\sigma_{ref}\in(50,500)$, and $\Delta\sigma\in(0,200)$ km s-1. Our log-likelihood function follows Appendix B of Bergamini et al. (2019).
@@
Each of the 81 possible permutations of these variables was run through pPXF with the population of 18 galaxies as described above to derive a new sample of stellar velocity dispersions, and the resulting spread provides us with the systematic error for our measurements. Fits with $\chi^{2}$ greater than 2.5 or an error on the stellar velocity dispersion greater than 75 km s-1 were removed from the sample for that run. We then re-fit the Faber-Jackson relation for each of the 81 realizations giving a range of values for our power-law fit parameters. In Table 1, we list the central 68% range on the velocity dispersions of the individual galaxies from these systematics runs.
@@ 4.1 Faber-Jackson Fit
Galaxy RA DEC F555w MARZ pPXF Velocity Systematics
Label Magnitude Redshift Redshift Dispersion (km s-1) Deviation (km s-1)
@@ 4.3 Significant Galaxies along the Line-of-Sight
As previously noted, some of the galaxies in our full redshift sample are not cluster members, but still lie along the line of sight to the lensed source and thus affect the lens modeling. Notably, galaxy A is the brightest in the field besides the central galaxy and one of the closest galaxies in projection to the central. Based on the pPXF analysis, this galaxy has a velocity dispersion of 236 km s-1, corroborating that the galaxy is of high mass. The spectrum and pPFX fit for this galaxy are shown in Figure 7. Along with A, we find four, fainter galaxies with similar redshifts, AF, AK, BC, and BD.
Figure 7: MUSE spectrum of galaxy A (black) overlaid with pPXF’s best fitting model (red) and residuals (green). The spectrum is shown from 3850Å to 4500Å and includes visible Calcium H & K and G-band absorption features.
@@ 4.4 X-ray Properties
From the Chandra analysis, we find that MACS0138 has an $r_{500}$ X-ray temperature of $6.7\pm 0.4$ keV and a soft band (0.5-2 keV) luminosity of $5.5\pm 0.1\times 10^{44}$ ergs s-1, with an $r_{500}$ radius of $1089^{+41}_{-38}$ kpc. The temperature is consistent for different aperture choices; for example, the $r_{2500}$ temperature is found to be $6.5^{+0.4}_{-0.3}$ keV at an $r_{2500}$ radius of $473^{+15}_{-13}$ kpc, and the core-cropped $r_{500}$ temperature is found to be $6.6^{+0.9}_{-0.8}$ keV. Based on the $r_{500}$ temperature and using the $M-T$ relation from Mantz et al. (2016), we estimate the mass of the cluster to be $M_{500}=5.2^{+1.5}_{-1.2}\times 10^{14}M_{\odot}$ where the uncertainties include both the X-ray temperature uncertainties and the uncertainties on the fit parameters in Mantz et al. (2016).
@@ 5 Discussion
For the Faber-Jackson relation, Bergamini et al. (2019) present results for three strong lensing clusters at similar redshifts to MACS0138 and find Faber-Jackson slopes of $\alpha=0.28\pm 0.02$, $0.27\pm 0.03$, and $0.27\pm 0.04$. Their results are very consistent with our result of $\alpha=0.26\pm 0.06(\mathrm{stat.})\pm 0.03(\mathrm{sys.})$, though we note that their reference velocity dispersions as slightly higher than ours ($\sim 300$ km s-1) and they include the BCG in their fits. Our measured intrinsic scatter is also consistent with the clusters in Bergamini et al. (2019) with the exception of Abell S1063 for which they measure a higher scatter.
In the final stages of preparation of this paper, the preprint Granata et al. (2024) was posted, which also fits the Faber-Jackson relation in MACS0138. Despite some differences in the two analyses, our results are completely consistent; they find $\alpha=0.25\pm 0.05$ and scatter $\Delta\sigma=25^{+4}_{-6}$ km s-1 at a similar reference velocity dispersion (206 km s-1 compared to our 223 km s-1). Furthermore, when fitting velocity dispersions of the same member galaxies, their analysis finds results largely consistent with those presented here. Their best fit velocity dispersions are all similar to ours and within the systematic and statistical uncertainties for every galaxy in common except AR and F for which we get larger velocity dispersion. For F, considering the uncertainties our value is within $\sim 2\sigma$ of what Granata et al. (2024) find; for AR our velocity dispersion places it relatively high compared to the Faber-Jackson relation while the Granata et al. (2024) value places it somewhat low. Differences in the two analyses include that we use a somewhat larger sample of galaxies (18 vs. 13) reflecting different quality cuts, they use a deeper MUSE cube obtained by their team, and they use F160W magnitudes compared to our use of F [^5] W. While they do investigate the affects of signal-to-noise on their velocity dispersion measurements, they do not employ the type of systematics test we do; our results indicate that their uncertainties may be underestimated. However, the overall consistency of the results of these independent works speaks to their robustness.
After the submission of this work, the preprint Acebron et al. (2025) was also posted using the same deeper MUSE data as Granata et al. (2024), and similarly conducts kinematic and X-ray analysis of cluster MACS0138. Their results for the Faber-Jackson slope excluding the BCG of $\alpha=0.26\pm 0.06$ is in good agreement with the findings presented here. Additionally, their findings for the X-ray temperature of 6-7 keV outside of 25 kpc, and thus their estimate of $M_{500}=5.7^{+3.4}_{-2.1}\times 10^{14}M_{\odot}$ are also consistent with our results (with an $r_{500}$ radius of 1070 kpc compared to 1089 kpc here)
@@ 6 Conclusion
The findings of this work are largely in agreement with similar independent works, such as Acebron et al. (2025) and Granata et al. (2024).